Given the following information, calculate the expected return and standard deviation for a portfolio that has 47 percent invested in Stock A, 15 percent in Stock B, and the balance in Stock C. (Do not round intermediate calculations. Enter your answers as a percent rounded to 2 decimal places.)

Returns
State of Economy Probability of State of Economy Stock A Stock B Stock C
Boom 0.70 10 % 23 % 26 %
Bust 0.30 13 0 −13
Answer #1

a). Expected Return = [Probability(i) * Return(i)]

Stock A's Expected Return = [0.7 * 10%] + [0.30 * 13%] = 7% + 3.9% = 10.9%

Stock B's Expected Return = [0.7 * 23%] + [0.30 * 0%] = 16.1% + 0% = 16.1%

Stock C's Expected Return = [0.7 * 26%] + [0.30 * 13%] = 18.2% - 3.9% = 14.3%

Portfolio's Expected Return = [0.47 * 10.9%] + [0.15 * 16.1%] + [(1 - 0.47 - 0.15) * 14.3%]

= 5.123% + 2.415% + 5.434% = 12.972%, or 12.97%

b). Portfolio's Standard Deviation = [{Probability(i) * (Expected Return - Return(i))2}]1/2

= [{0.47 * (12.972% - 10.9%)2} + {0.15 * (12.972% - 16.1%)2} + {0.38 * (12.972% - 14.3%)2}]1/2

= [2.0178%2 + 1.4677%2 + 0.6702%2]1/2 = [4.1556%2]1/2 = 2.04%